Grade 12th passPhysical Chemistry

please who can i calculate the amount of atom in 4.0 g of carbon? (c=12)

johnson Aniekan
8 Years agoGrade 12th pass

3 Answers

Ravneet
ApprovedApproved Tutor Answer8 Years ago
the number of moles = given mass/molar mass 
 
step 1: calculate the number of moles, that is 4/12 = 1/3
step 2: to calculate the number of atoms multiply by Avogadro number because  a mole of substance contains Avogadro number of entities.
so answer is 1/3 X Na
MonishGayen
8 Years ago
no. of moles of carbon =(given mass)/(molar mass)
=(4/12)NA​ =(1/3)*6*1023 =2*1023​(approximately)
lets hope it helps you
 
Abhinav
8 Years ago
4/12=1/3, therefore there is 1/3 moles of carbon
By unitary method 
No of atoms in 1/3 moles is 1/3*avagadro number